LOFAR Newsletters June 2016

Distributed on 1 July 2016

------------------

 

Array status:

----------------------

-       38 stations operational in the Netherlands: 24 core stations and 14 remote stations

-       12 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609, PL610, PL611, PL612

-       A new LOFAR station (IE613) will be built on the grounds of Birr Castle, located centrally in Ireland. Expected completion: first half 2017.

-       Oscillating tiles were detected during BF observations at a few stations and were deactivated.

-       The maintenance season has started and will continue till the weather conditions will allow.

-       The overview of non-operational antenna elements for LBA and HBA is available here.

-       Station calibration:

  • New calibration tables were installed for LBA OUTER in the second half of May. The calibration of the LBA SPARSE modes is still on-going.
  • It has been found challenging to calibrate the inner part of the array in SPARSE modes. Further data collections in these modes are foreseen to be needed.
  • New calibration campaigns will follow the station maintenance.

 

Observing System Status:

------------------------------------ 

-       The observing system experienced instability during the period April - June 2016. Problems were associated with the processing and LTA (ingest) systems and COBALT nodes. These caused delays in the exploitation of the Cycle 5 & 6 observing programmes and in digesting data out of CEP2.

 

-       The Radio observatory has discovered that a transient issue with the RSP drivers and the beam server has been impacting station sensitivities, potentially affecting correlated as well as beam-formed observations. The conditions that caused the issue have always been present in the system, but the issue manifested more prominently last January. The problem affected the first few minutes of observations performed in multiple scans (i.e. calibrator-target-calibrator setup or shifting between targets), where the gap between observations was less than three minutes. A fix has been rolled-out to the stations and the issue is now resolved. Further details will be circulated in the next few days.

 

Software Development Status (J. Annyas):

------------------------------------------

-       CEP4 & Resource Assigner 1.0 were tested by Software Support and have been recently handed over to Science Support for commissioning.

-       The new LCUs for all station have been ordered and the new software+hardware combination is being tested.

-       The NorthStar proposal submission tool is undergoing debugging in view of the Cycle 7 proposal call.

 

CITT Status (T. J. Dijkema & E. Orru')

--------------------------------------------------------------------

-       The calibration & Imaging Tiger Team is preparing for the next Imaging Busy Week (4-8 July). The focus of this week is on improving the prefactor (direction independent) and factor (direction dependent) imaging pipelines.

-       DPPP gaincal has had various updates to make calibration with DPPP more convenient. A new feature to fit for TEC will be tested during the busy week.

-       Developments on the new fast gridder are ongoing. A paper about the gridder will be submitted soon.

 

MSSS (J. Broderick)

Progress toward the final MSSS-HBA catalog continues, with an enhanced flux scale correction methodology underway. A direct empirical determination of flux scale corrections gave sensible improvements in a broad declination range, but not over the full survey area. Improvements that are applicable for the full Northern sky are now being developed through a constrained bootstrapping technique. In addition, fortnightly videocons have restarted to provide a forum for commissioning and science updates; if you are interested in participating, please email Jess Broderick (broderick [at] astron [dot] nl)

 

Observing Programmes

--------------------------------------------------------------------

-       Cycle 6 Allocations are summarised here. The observing schedule can be fund here.

-       During the Cycle 6 PC meeting, processing time proved to be a limiting factor for the next Cycle. An over-allocation of processing time was agreed, but with the associated observing time being given with low priority.  In the schedule, observations are labelled as priority "A" or "B". "B" observations will only be scheduled if processing and data ingest can be accommodated at the time.  If not, they will be moved to later in the Cycle, but will not carry on to the next Cycle.

-       Cycle 6 observing programme: 18% complete.

-       Cycle 5 observing programme: 92% complete. The rest is being observed with second priority during Cycle 6.

 

Documentation

--------------------------------------------------------------------

-       A collection of LOFAR slides that can be used when preparing LOFAR-related presentations is available here.

-       LOFAR science highlights are collected here and in ppt form here.

-       The list of LOFAR papers is available here.

-       Documentation about LOFAR - including the presentations from the LOFAR family meeting 2016 - is available here.

 

CEP news:

----------------

-       CEP4

  • The CEP4 cluster is being commissioned.

-       CEP2:

  • Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.
  • During the past two months, a few locus nodes experienced issues and were temporarily taken out of production.

-       CEP3:

  • Cluster info and schedule available here.
  • CEP3: following an analysis of the usage of CEP3 and requests for extensions received during the past year, the policy to access to the cluster has been slightly modified. Default period extended from 4 weeks to 8 weeks.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

Note: the following events are now marked on an online calendar that is available here.

-       25th Imaging Busy Week: 4-8 July

-       Cycle 7 proposal call: July 2016

-       LOFAR Data Processing School: 5-9 September 2016

-       Next LSM's : 06/07, 20/07, 17/08, 31/08

(note: all presentations given at the LSM are available here).

-       Next Stop days: 12/07, 02/08, 30/08

-       Next software roll outs: 04/07

-       Next LOFAR bulletin: August 2016.

 

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON